OpenStudy (anonymous):

A sample of radioactive material contains 1.00e^15 atoms and has an activity of 6.00e^11 Bq. What is the half life?

OpenStudy (anonymous):

\[A=A_0 e^{rt}\] \[.5= e^{(-6.00 e^{11})t}\] natural log both side and solve for t \[1.92946*10^{-6}\] I would take it with grain of salt,

OpenStudy (anonymous):

i have 1.16e^3 s..may be you have minutes.let me convert

OpenStudy (anonymous):

no i mean the answer in the book(sorry)

OpenStudy (anonymous):

i think it is e^(negative power)

OpenStudy (anonymous):

but yh u have -ive already

OpenStudy (anonymous):

i meant r= -r

OpenStudy (anonymous):

can you recheck all the numbers

OpenStudy (anonymous):

ok lemme see

OpenStudy (anonymous):

I have the same formula but instead of r the book uses lambda. and it is negative

OpenStudy (anonymous):

so did the book used \[\left(6* 10^{-4}\right)\]for r?

OpenStudy (anonymous):

Ok so what i have here is R=Ro e^-(lambda)t

OpenStudy (anonymous):

What did they plug in for lambda

OpenStudy (anonymous):

lambda= Ro/No

OpenStudy (anonymous):

No= \[5.98741\times 10^8\] lambda=6* 10^-4 \[\frac{1}{2}\text{==}E^{-\left(6\ 10^{-4}\right)t}\]

OpenStudy (anonymous):

wow nice!!

OpenStudy (anonymous):

thanks a lot:)

OpenStudy (anonymous):

not really sure what N0 means , but it works

OpenStudy (anonymous):

No= number of undecayed atoms

OpenStudy (anonymous):

now back to my own physics